T. In a population, 60% of the people have brown hair, 40% have brown e\es and 30% have brown hair and brown eyes.

a. What percent of those who have brown hair will have brown eyes?

b. According to a recent mortality table, the probability that a 35-\ear-old US citizen will live to age 65 is 0.725. John and Jim, two 35-year-old Americans who are not relatives, L. What is the probability that neither man lives to age 65? LL. What is the probability that both live to age 65?
Answer #1

P( people having brown hair) = 0.6

P( people having brown eyes) = 0.4

P( people having brown hair and brown eyes ) = 0.3

A) Percent of  those who have brown hair will have brown eyes = ( percent having both brown hair and brwon eyes / percent having brown hair) * 100

= 0.3/0.6*100

= 50%

B)   Probability (A 35 year-old US citizen will live to age 65) =  0.725

P( neither man lives to age 65) = P( john dies before reaching age 65) * P( jim dies before reaching age 65)

= (1 - 0.725) * (1 - 0.725)

= 0.075625

P( both man lives to age 65) = P( john lives to age 65) * P( jim lives to age 65)

= 0.725 * 0.725

= 0.525625